Operations Report Apr 21st

Crew 245 Operations Report 21APR2021

SOL: 11

Name of person filing report: Shravan Hariharan

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator: Working nominally

Hours run: 7

From what time last night: 9:30 pm, SOC 60%

To what time this morning: 4:30 am, SOC 100%

List any additional daytime hours when the generator was run: N/A

Solar— SOC 65% at 7:48 pm (Before generator is run at night)

Notes on power system: For the last several sols, the freezers in the Science Dome have been running in order to preserve science samples. Due to this, the station power consumption has been significantly higher than it previously was, resulting in the generator being turned on earlier (as early as 7:00 pm on colder days, and 9:00 pm on warmer days) once the SOC reaches below 70%.

Diesel Reading – 50%

Station Propane Reading – 70%

Water (loft tank): 25 gallons

Water Meter: 150665.5 units

Water (static tank): 135 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Spirit rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Opportunity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: no, in town

Curiosity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Perseverance rover used: No

Hours: 204.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Left tail-light on Perseverance rover is not functioning – noticed on EVA.

ATV’s Used: 350.2

Reason for use: EVA

Oil Added? No

ATV Fuel Used: 0.5 gallon (estimated)

# Hours the ATVs were used today: 4 (but 0.5 hour of active transit)

Notes on ATVs: Mission Director will monitor and refuel ATVs for duration of mission.

HabCar used and why, where?: Nothing to report

CrewCar used and why, where?: Nothing to report

General notes and comments: Nothing to report

Summary of internet: Internet is working nominally. Crew is rationing internet usage during the day to ensure that there is sufficient bandwidth for Mission Support Comms window.

Summary of suits and radios: All radios nominal, all suits currently charging. Suits 7, 9, and 11 used on EVA today, and airflow reduced for all suits approximately 3 hours into EVA. No further EVAs planned with these suits, but Crew Engineer suspects issue is with the chosen charging outlet – all three suits were charging from same power strip. Crew Engineer will test these suits tosol to determine root cause of issue and any potential fixes.

Summary of Hab operations: Flapping side panel on outside of Hab. Station Director secured panel with 2×4, so it is no longer making noise

Summary of GreenHab operations: Being used to host Crew Botanist regolith feasibility experiment, but no active operations.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report